OpenStudy (anonymous):

if 1.2 kilograms of rust forms on a bridge in 5 days, what should be the rate of reaction in grams per hour?

OpenStudy (anonymous):

1.2Kg/5day*1000g/Kg*1day/24hours = 10.0 grams/hour

OpenStudy (unklerhaukus):

\[{1.2kg \over 5day} \times {1000g\over kg}\times{1day\over 24hr} = 10.0 {g \over hr}\] iv have just rewritten Jetzt answer to make it more readable